# Provable Emergence of Deep Neural Collapse and Low-Rank Bias in $L^{2}$-Regularized Nonlinear Networks

### v2

@@ 4 Vanishing $\operatorname{\mathrm{TCV}}$: global optimality of deep-neural collapse
The proof of Theorem˜4.3 can be found in Section˜B.1. The main takeaway is that, from almost every interpolating initial condition (aside from a set of measure zero), one can construct a continuous path in parameter space along which the loss decreases (up to a finite set of downward jumps) and the network converges to a global minimum characterized by deep neural collapse. We highlight that in any numerical method, one looks at a discretized version of the path $\gamma$, so the “probability" of seeing discontinuity points is zero, no matter how small the learning rate. In particular, no matter how small the stepsize, there always exists a sequence of points that decreases the loss and reaches $\operatorname{\mathrm{DNC1}}$ configurations. This result is exactly formalized in Corollary˜B.4 and Remark˜B.5. In particular, a consequence of Theorem˜4.3 is that $\inf_{\gamma\in\Gamma}\sup_{t\in[0,1]}\mathscr{L}(\gamma(t))\leq\mathscr{L}(Z(0))$ where $\Gamma:=\{\gamma\in\mathcal{C}^{0}([0,1];\mathcal{S})\,|\,:\gamma(0)=Z(0),\gamma(1)\,\text{satisfies}\,\,\operatorname{\mathrm{DNC1}}\,\,\forall l=1,\dots,L\}$.
@@
A second way to interpret Theorem˜4.3 and Corollary˜B.4 is that the loss landscape of $\mathscr{L}$, despite being highly nonconvex, is well behaved in the sense that there are essentially no loss barriers from almost all interpolating points to global minima. This result, being independent of the specific temporal dynamics imposed during training, provides theoretical support for the empirical “universality” of neural collapse observed in classification tasks when using sufficiently wide architectures Poggio_hidden_nc and possibly gives an hint on why it is observed towards the end of training papyan2020prevalence ; Poggio_hidden_nc (i.e., in the interpolating phase, where the landscape is well behaved).
@@
Moreover, the phenomenon is “unidirectional”, progressing from high to low rank as demonstrated in the proof, which is reminiscent of the behaviour observed in the deep linear case in wang2024implicit .
@@
Finally, we point out that these results are recovered as the singular limit in the regime where an arbitrarily small penalization term for the first layer $\|W_{1}\|^{2}=\|\sigma_{1}^{-1}(Z^{1})X^{+}\|^{2}$ is included. A natural question arises: are the highly structured solutions given by Theorem˜4.2 preserved when $\|W_{1}\|^{2}$ is included? To answer this question about stability, we consider the perturbed loss
@@
$\mathscr{L}_{\lambda}(Z)=\begin{cases}+\infty,\quad\text{if}\,\,\ Z^{L}\neq Y,\\ \frac{1}{2}\sum_{l=1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{1}{\lambda}\|\sigma_{1}^{-1}(Z^{1})X^{+}\|^{2},\quad\text{otherwise},\end{cases}$ (12)
@@
defined on the set $\mathcal{S}$, see (4.1), and show that, as $\lambda\to\infty$, the global minima of $\mathscr{L}_{\lambda}$ converge to global minima of $\mathscr{L}$. Note that varying the value of $\lambda$ is equivalent to rescaling the dataset $X$ by a factor $\lambda$.
@@
The functionals $\mathscr{L}_{\lambda}:\mathcal{S}\to\mathbb{R}$ converge to $\mathscr{L}:\mathcal{S}\to\mathbb{R}$ as $\lambda\to+\infty$, in the $\Gamma$-convergence sense.
Moreover, as a classical consequence of $\Gamma$-convergence, if $Z^{*}\in\mathcal{S}$ is an accumulation point for a sequence of minimizers $Z_{n}\in\arg\min_{Z\in\mathcal{S}}\mathscr{L}_{\lambda_{n}}(Z)$ with $\lambda_{n}\to+\infty$, then $\mathscr{L}(Z^{*})=\min_{Z\in\mathcal{S}}\mathscr{L}(Z)$. The proof is included in Section˜B.1. We also highlight here that the existence of an accumulation point $Z^{*}$ is not immediate, since $\mathcal{S}$ is not compact.
The interpretation of Theorem˜4.4 is twofold. First, the optimality of $\operatorname{\mathrm{DNC1}}$ configurations in Theorems˜4.2 and C.2 are stable with respect to the presence of regularization on $\|W_{1}\|^{2}$, therefore, we expect configurations close to $\operatorname{\mathrm{DNC1}}$ for all layers even when the first weight decay is present. Moreover, the regime in which one may expect configurations to be close to $\operatorname{\mathrm{DNC1}}$ ones with the first weight decay included is when $\|X\|\gg\|Y\|.$ Alternatively, another possible interpretation of $\lambda$ in Theorem˜4.4 is as the depth of the network. In fact, the minimizers of Equation˜12 and those of Equation˜12 divided by $L$ are the same. The first term, being bounded on the minimizers, is vanishing for $L\to+\infty$. Therefore, Theorem˜4.4 is essentially telling that one should expect the minimizer structure as the one predicted in Theorem˜4.2 for large depth networks. Furthermore, we remark that the regularization term due to any finite number of initial layers vanishes in the large depth limit, making the results of Theorem˜4.2 hold also for less restrictive conditions on the widths, but from the first layer $l^{*}$ which can represent a large enough region of the full space $\mathbb{R}^{n_{l^{*}}\times N}$ (the one containing the minimizers predicted in Theorem˜4.2). A more precise discussion of this can be found in Remark˜B.6. An equivalent stability result holds even for residual networks, as discussed in Remark˜C.3.
@@ 5 Experimental evaluation
5.1 Qualitative behaviour of Theorem˜3.4
In this section, we show in a simple setting the qualitative dependency of $\min_{Z\in\cup_{r=1}^{K}\mathcal{M}_{r}}\|W^{*}_{l}-Z\|$ on the regularization parameter $\lambda$ and on the TCV of the dataset. In particular, fixed a set of vectors $\mu_{1},\dots,\mu_{K}$, we consider a random Gaussian mixture dataset $X\in\mathbb{R}^{n\times N}$ with $x_{i}\sim N(\mu_{y_{i}},\sigma^{2})$, and a $Y\in\mathbb{R}^{K\times N}$ one-hot encoding matrix with $K=10$ classes. In this way, we know that $\operatorname{\mathrm{TCV}}_{K,0}\sim O(\sigma^{2})$, and we expect, from Theorem˜3.4, the ranks to approach $K=10$ as $\sigma^{2}\to 0$. Then, for each $(\lambda,\sigma)$ in a grid of values, we sampled $X,Y$ according to the distribution described above, and we trained a four-layer fully connected neural network with the loss defined in Equation˜3. In Figure˜1 we report the results of this experiment. We can observe that consistently, for each $\lambda$ in our grid, we have that all matrices are essentially of rank $K$ for $\sigma$ small enough. Moreover, we also observe the decreasing behavior with respect to $\lambda$ consistently with Theorem˜3.4.
@@
5.2 Convergence to the optimal solution of Theorem˜4.3
We start by illustrating and numerically validating what we claimed in Theorems˜4.3 and 4.4. In particular, by setting $\delta=1/\lambda$ in (12) we can show that for $\delta\to 0$, the singular values of all the intermediate post-activation features are the ones described in Theorem˜4.2. To set up the experiment, we subsampled the MNIST dataset deng2012mnist with $N=200$ examples, and we considered a $5$ layer feedforward network with all $\sigma_{l}=\tanh$ activations and widths equal to $n_{1}=\dots=n_{4}=N=200,n_{5}=K=10$. With $\delta$, the loss becomes $\mathcal{L}_{\lambda,\delta}=\frac{1}{2}\|f_{\mathit{\Theta}}(X)-Y\|^{2}+\frac{\lambda}{2}\sum_{l=2}^{L}\|W_{l}\|^{2}+\frac{\lambda\delta}{2}\|W_{1}\|^{2}+\frac{\lambda^{\prime}}{2}\sum_{l=1}^{L}\|Z^{l}\|^{2}$, where the last regularization term can be thought of as a weaker way of imposing the second set of constraints in Theorem˜4.3. In particular, for this experiment we considered $\lambda=\lambda^{\prime}=10^{-5}$.
We trained the network until convergence (way past interpolation, with a final data loss lower than $10^{-6}$ in all examples) for different values of $\delta$, and we compared the actual singular values of the feature matrices after training with the ones predicted by Theorem˜4.2. The results are shown in Figure˜2. As we can notice, the singular values are converging to the ones predicted by Theorem˜4.2, and the convergence with respect to $\delta$ seems to be fast. In fact, already for moderate sizes of $\delta$ (e.g., $\delta=2$), the singular values are close to the predicted ones. Additional results are included in Figure˜4.
$\delta=10$ $\delta=5$ $\delta=2$ $\delta=0.01$
@@
Figure 2: Singular values of trained networks (full line) against those predicted by Theorem˜4.2 (dots) for different values of $\delta=\frac{1}{\lambda}\in\{10,5,2,0.1\}$. The singular values are converging to the predicted ones as $\delta\to 0$, as shown by Theorem˜4.4 in combination with Theorem˜4.2.
5.3 Convergence to $\operatorname{\mathrm{DNC1}}$ configurations
In this experiment, we will showcase the convergence of $\operatorname{\mathrm{TCV}}$ to zero of all intermediate layers in the setting of Theorem˜4.2. The experimental setting is the same as in Section˜5.2, but with $L=10$ layers and $\delta=0$ fixed. In Figure˜3 (rightmost figure), we plot for a single run the $\operatorname{\mathrm{TCV}}$ of all intermediate post-activation embeddings. As we can observe from this first plot, the $\operatorname{\mathrm{TCV}}$ of all intermediate embeddings goes essentially to machine precision as expected from Theorems˜4.2 and 4.3. In Figure˜3 (first three figures from the left), we instead plot the average $\operatorname{\mathrm{TCV}}$ across all intermediate layers, for $100$ different random unitary initializations, to avoid slow convergence caused by the depth. As we can see, the $\operatorname{\mathrm{TCV}}$ consistently goes to zero in all runs. Regarding the numerical ranks (up to $10^{-5}$ relative error), we observe that they are all close to the optimal ones ($K=10$).
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: First three plots from the left: average of different quantities on all intermediate layers and for $100$ different random initializations. Shadows represent the area included between the minimal and maximal values observed. Each optimization step represents 100 epochs. Right figure: TCV of each single layer during training.
@@ Appendix A Useful definitions and results
In this section, we will introduce some useful definitions and fact which we will need to prove the results in Appendix˜B. We start recalling the definition and some basic facts about $\Gamma$-convergence of functionals (see [6]).
@@
[6] Let $F_{n}\to F$ in the sense of $\Gamma$-convergence and let $S$ be sequentially compact. Then, for every sequence $x_{n}\in\arg\min_{y\in S}F_{n}(y)$ and every limit point $x$ of $(x_{n})$, we have
@@
The next lemma discusses the zeros of single-variable analytic functions, showing that their cardinality is always finite on compact sets, which we will need later.
Lemma A.3 (Finite number of zero of real analytic functions on compact intervals, [19]).
@@
Let $X,Y$ be two finite-dimensional normed spaces, $f:X\times Y\to\mathbb{R}$ and $g:X\times Y\to\mathbb{R}^{m}$ be continuous functions, and $\Omega(x)=\{y\in Y\,|g(x,y)=0\,\}$. Moreover, assume that $f(x,y)$ is coercive with respect to $y$, uniformly in $x$. Then, the function
@@
We will use this last result in Section˜B.1 to prove the lower semicontinuity of $F(A,B)=\|AB^{+}\|^{2}$. In fact, by defining
@@
we can define g(x,y)=WBAg(x,y)=WB-A, $f(x,y)=\|W\|^{2}$ and
$F(x)=\inf_{y\in\Omega(x)}f(x,y),$
recalling that $\Omega(x)=\{y=W\;|\;g(x,y)=WB-A=0\}$ and using a matrix version of [10, Section 5.5.1]. In particular we are exactly in the setting of Proposition˜A.4, proving lower semicontinuity of $F(A,B)=\|AB^{+}\|^{2}$.
@@
$f(B_{k})$ $=\|AV_{k}^{1}\Sigma_{k,1}^{-1}U_{k,1}^{T}+AV_{k,2}\Sigma_{k,2}^{-1}U_{k,2}^{T}\|\geq\|AV_{k,2}\Sigma_{k,2}^{-1}\|-\|AV_{k,1}\Sigma_{k,1}^{-1}\|$ (13)
@@
The main step of the proof is to show that the gradient with respect to a single data point produces matrices of rank at most 1. One then concludes by observing that the sum of $K$ terms of rank at most $1$ has rank at most $K$. Given the structure of the loss function that depends only on the output of the model $f(\mathit{\Theta},x)$, if we use the chain rule and the recursive definition of $f(\mathit{\Theta},x)$ in (1), we obtain
@@
$\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},z_{i}),y_{i}\big{)}=$ $\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}$ (14)
@@
$R_{i}^{k}(\mathit{\Theta}):=\frac{1}{2}\operatorname{\nabla}_{(z,y)}^{2}\nabla_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\alpha_{i}^{k})\beta_{i}^{k}\big{)}[(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k}),(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k})],$ (15)
for some $(\alpha_{i}^{k},\beta_{i}^{k})$ between $(z_{i},y_{i})$ and $(\bar{z}_{k},\bar{y}_{k})$. Therefore, using the linearity of $\operatorname{\nabla}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k})\bar{y}_{k}\big{)}[\cdot]$, we have
@@
$M_{\operatorname{\ell},j}(\mathit{\Theta})=\sup_{k=1,\dots,K}\sup_{i\in\mathcal{C}_{k}}\|\operatorname{\nabla}^{2}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\alpha_{i}^{k}),\beta_{i}^{k}\big{)}\|\leq\sup_{k=1,\dots,K}\sup_{(\alpha,\beta)\in\mathcal{H}_{k}^{j}}\|\operatorname{\nabla}^{2}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\alpha),\beta\big{)}\|,$ (16)
@@
A.3 Proof of Theorem˜3.4
Observe that from (4), at the stationary point we have
@@
Now, Lemma 3.3 allows us to control the distance of the gradient in the right-hand side of (17) from $\mathcal{M}_{K}^{d_{l-1},d_{l}}$. In particular, given $\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}$, for any $j=1,\dots,l-1$ we have

### v3

@@ 4.1 Characterizing Optimal Representations: Natural Architectural Constraints
The full proof of Theorem˜4.4 can be found in Section˜B.2. The main takeaway is that, from almost every interpolating initial condition (aside from a set of measure zero), one can construct a continuous path in parameter space along which the loss decreases (up to a finite set of downward jumps) and the network converges to a global minimum characterized by deep neural collapse. In particular, Theorem˜4.4 implies that
$\inf_{\gamma\in\Gamma}\sup_{t\in[0,1]}\mathscr{L}(\gamma(t))\leq\mathscr{L}(Z(0)),$
where $\Gamma:=\{\gamma\in\mathcal{C}^{0}([0,1];\mathcal{S})\,|\,:\gamma(0)=Z(0),\gamma(1)\,\text{satisfies}\,\,\operatorname{\mathrm{DNC1}}\,\,\forall l=1,\dots,L\}$.
We highlight that in any numerical method, one looks at a discretized version of the path $\gamma$, so the “probability" of seeing discontinuity points is zero, no matter how small the learning rate. In particular, no matter how small the stepsize, there always exists a sequence of points that decreases the loss and reaches $\operatorname{\mathrm{DNC1}}$ configurations. This result is exactly formalized in the following corollary, together with a more detailed explanation in Remark˜B.6:
Corollary 4.5 (No loss barrier for any finite stepsize).
In the setting of Theorem˜4.4, for almost every $Z(0)\in\mathcal{S}$ there exists a sequence of curves $\gamma_{k}:[0,1]\to\mathcal{S}$ uniformly converging to the continuous curve $\gamma$ of Theorem˜4.4 such that the following hold for all $k\geq 1$:
  • $\gamma_{k}$ is piecewise constant in a set of disjoint intervals $\{I_{j,k}:=[T_{j,k},T_{j+1,k})\}_{j=0,\dots,k-1}$ whose union is $[0,1]$, and with $0<T_{j+1,k}-T_{j,k}<\frac{1}{k}$ for all $j=0,\dots,k-1$;
  • $\gamma_{k}(0)=Z(0)$, $\gamma_{k}(1)\in\underset{Z\in\mathcal{S}}{\arg\min}\,\mathscr{L}(Z)$
  • $\mathscr{L}$ is decreasing, i.e., $\mathscr{L}(\gamma_{k}(0))\geq\mathscr{L}(\gamma_{k}(T_{1,k})))\geq\mathscr{L}(\gamma_{k}(T_{2,k})))\geq\dots\geq\mathscr{L}(\gamma_{k}(T_{k,k})))\geq\mathscr{L}(\gamma_{k}(1)).$
The idea of the proof Corollary˜4.5 relies essentially on a discretization of the path from Theorem˜4.4, and can be found in Section˜B.4.
@@
A second way to interpret Theorem˜4.4 and Corollary˜4.5 is that the loss landscape of $\mathscr{L}$, despite being highly nonconvex, is well behaved in the sense that there are essentially no loss barriers from almost all interpolating points to global minima. This result, being independent of the specific temporal dynamics imposed during training, provides theoretical support for the empirical “universality” of neural collapse observed in classification tasks when using sufficiently wide architectures (Rangamani et al., 2023a) and possibly gives an hint on why it is observed towards the end of training (Papyan et al., 2020; Rangamani et al., 2023a) (i.e., in the interpolating phase, where the landscape is well behaved).
@@
Moreover, the phenomenon is “unidirectional”, progressing from high to low rank as demonstrated in the proof, which is reminiscent of the behaviour observed in the deep linear case in (Wang and Jacot, 2024).
@@
Finally, we point out that these results are recovered as the singular limit in the regime where an arbitrarily small penalization term for the first layer $\|W_{2}\|^{2}=\|\sigma_{2}^{-1}(Z^{2})Z^{1,+}\|^{2}$ is included. A natural question arises: are the highly structured solutions given by Theorem˜4.2 preserved when $\|W_{1}\|^{2}$ is included? To answer this question about stability, we consider the following perturbed loss, where $\phi$ is a generic nonnegative lower-semicontinuous function defined on the set $\mathcal{S}$ (Definition˜4.1),
@@
$\mathscr{L}_{\lambda}(Z)=\begin{cases}+\infty,\quad\text{if}\,\,\ Z^{L}\neq Y,\\ \frac{1}{2}\sum_{l=1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{1}{\lambda}\phi(Z),\quad\text{otherwise},\end{cases}$ (13)
@@
and show that, as $\lambda\to\infty$, the global minima of $\mathscr{L}_{\lambda}$ converge to global minima of $\mathscr{L}$. Note that varying the value of $\lambda$ is equivalent to rescaling the dataset $X$ by a factor $\lambda$.
@@
Assume $\phi:\mathcal{S}\to\mathbb{R}_{+}$ is lower semicontinuous. Then, the family of functionals $\mathscr{L}_{\lambda}:\mathcal{S}\to\mathbb{R}$ converge to $\mathscr{L}:\mathcal{S}\to\mathbb{R}$ as $\lambda\to+\infty$, in the $\Gamma$-convergence sense.
Moreover, as a classical consequence of $\Gamma$-convergence, if $Z^{*}\in\mathcal{S}$ is an accumulation point for a sequence of minimizers $Z_{n}\in\arg\min_{Z\in\mathcal{S}}\mathscr{L}_{\lambda_{n}}(Z)$ with $\lambda_{n}\to+\infty$, then $\mathscr{L}(Z^{*})=\min_{Z\in\mathcal{S}}\mathscr{L}(Z)$. The proof is included in Section˜B.2. We also highlight here that the existence of an accumulation point $Z^{*}$ is not immediate, since $\mathcal{S}$ is not compact. We finally remark that an equivalent stability result holds even for residual networks, as discussed in Remark˜C.3.
Remark 4.7.
Theorem˜4.6 can be interpreted from two complementary perspectives. First, the optimality of $\operatorname{\mathrm{DNC1}}$ configurations established in Theorems˜4.2 and C.2 remains stable under the inclusion of a backbone regularization term. Consequently, one expects the configurations across all layers, up to the first representative layer, to remain close to $\operatorname{\mathrm{DNC1}}$. In particular, such proximity to $\operatorname{\mathrm{DNC1}}$ configurations, even when the first-layer weight decay is incorporated, is expected to occur in the regime $\|X\|\gg\|Y\|$.
Alternatively, the parameter $\lambda$ in Theorem˜4.6 can also be interpreted as representing the network depth. To formalize this interpretation, define
$\mathscr{L}^{[l_{1},l_{2}]}_{L}(Z):=\begin{cases}+\infty,&\text{if }Z^{L}\neq Y,\\[3.0pt] \frac{1}{L}\displaystyle\sum_{l=l_{1}}^{l_{2}}\phi_{l+1}(Z^{l+1},Z^{l}),&\text{otherwise.}\end{cases}$
For $Z^{L}=Y$, consider the difference
$\mathscr{L}^{[0,L-1]}_{L}-\mathscr{L}^{[l^{*},L-1]}_{L}=\mathscr{L}^{[0,l^{*}-1]}_{L}=\frac{1}{L}\sum_{l=0}^{l^{*}-1}\phi_{l+1}(Z^{l+1},Z^{l})\underset{\Gamma}{\longrightarrow}0,\quad\text{as }L\to+\infty.$
Hence, in the large-depth regime, even in the absence of a pyramidal backbone structure, one expects the minimizers to exactly satisfy the $\operatorname{\mathrm{DNC1}}$ condition from the first layer $l^{*}$ for which the set of representable features $Z^{l^{*}}$ contains the minimizer predicted by Theorem˜4.2, with the special case of the UFM, where the set of representable matrices $Z^{l^{*}}$ coincides with $\mathbb{R}^{n_{l^{*}}\times N}$.
This observation elucidates that, in general, for sufficiently deep architectures, the structure of the minimizers asymptotically approaches that predicted in Theorem˜4.2, where $Z^{2}$ is effectively replaced by the minimal layer index $l^{*}$ such that $Z^{l^{*}}$ is representative enough.
@@ 5 Experimental Evaluation
5.1 Qualitative Behaviour of Theorem˜3.5
In this section, we show in a simple setting the qualitative dependency of $\min_{Z\in\cup_{r=1}^{K}\mathcal{M}_{r}}\|W^{*}_{l}-Z\|$ on the regularization parameter $\lambda$ and on the TCV of the dataset. In particular, fixed a set of vectors $\mu_{1},\dots,\mu_{K}$, we consider a random Gaussian mixture dataset $X\in\mathbb{R}^{n\times N}$ with $x_{i}\sim N(\mu_{y_{i}},\sigma^{2})$, and a $Y\in\mathbb{R}^{K\times N}$ one-hot encoding matrix with $K=10$ classes. In this way, we know that $\operatorname{\mathrm{TCV}}_{K,0}\sim O(\sigma^{2})$, and we expect, from Theorem˜3.5, the ranks to approach $K=10$ as $\sigma^{2}\to 0$. Then, for each $(\lambda,\sigma)$ in a grid of values, we sampled $X,Y$ according to the distribution described above, and we trained a four-layer fully connected neural network with the loss defined in Equation˜3. In Figure˜1 we report the results of this experiment. We can observe that consistently, for each $\lambda$ in our grid, we have that all matrices are essentially of rank $K$ for $\sigma$ small enough. Moreover, we also observe the decreasing behavior with respect to $\lambda$ consistently with Theorem˜3.5.
@@
5.2 Convergence to the Optimal Solution of Theorem˜4.4
We start by illustrating and numerically validating what we claimed in Theorems˜4.4 and 4.6. In particular, by setting $\delta=1/\lambda$ in equation 13 we can show that for $\delta\to 0$, the singular values of all the intermediate post-activation features are the ones described in Theorem˜4.2. To set up the experiment, we subsampled the MNIST dataset (Deng, 2012) with $N=200$ examples, and we considered a $5$ layer feedforward network with all $\sigma_{l}=\tanh$ activations and widths equal to $n_{1}=\dots=n_{4}=N=200,n_{5}=K=10$. By including $\delta$, the loss becomes $\mathcal{L}_{\lambda,\delta}=\frac{1}{2}\|f_{\mathit{\Theta}}(X)-Y\|^{2}+\frac{\lambda}{2}\sum_{l=2}^{L}\|W_{l}\|^{2}+\frac{\lambda\delta}{2}\|W_{1}\|^{2}+\frac{\lambda^{\prime}}{2}\sum_{l=1}^{L}\|Z^{l}\|^{2}$, where the last regularization term can be thought of as a weaker way of imposing the second set of constraints in Theorem˜4.4. In particular, for this experiment we considered $\lambda=\lambda^{\prime}=10^{-5}$.
We trained the network until convergence (way past interpolation, with a final data loss lower than $10^{-6}$ in all examples) for different values of $\delta$, and we compared the actual singular values of the feature matrices after training with the ones predicted by Theorem˜4.2. The results are shown in Figure˜2. As we can notice, the singular values are converging to the ones predicted by Theorem˜4.2, and the convergence with respect to $\delta$ seems to be fast. In fact, already for moderate sizes of $\delta$ (e.g., $\delta=2$), the singular values are close to the predicted ones.
$\delta=10$ $\delta=2$ $\delta=1$
Refer to caption Refer to caption Refer to caption
$\delta=10^{-1}$ $\delta=10^{-2}$ $\delta=0$
Refer to caption Refer to caption Refer to caption
Figure 2: Singular values of trained networks (full line) against those predicted by Theorem˜4.2 (dots) for different values of $\delta=\frac{1}{\lambda}\in\{10,2,1,0.1,0.01,0\}$. The singular values are converging to the predicted ones as $\delta\to 0$, as shown by Theorem˜4.6 in combination with Theorem˜4.2.
5.3 Convergence to $\operatorname{\mathrm{DNC1}}$ Configurations
In this experiment, we will showcase the convergence of $\operatorname{\mathrm{TCV}}$ to zero of all intermediate layers in the setting of Theorem˜4.2. The experimental setting is the same as in Section˜5.2, but with $L=10$ layers and $\delta=0$ fixed. In Figure˜3 (rightmost figure), we plot for a single run the $\operatorname{\mathrm{TCV}}$ of all intermediate post-activation embeddings. As we can observe from this first plot, the $\operatorname{\mathrm{TCV}}$ of all intermediate embeddings goes essentially to machine precision as expected from Theorems˜4.2 and 4.4. In Figure˜3 (first three figures from the left), we instead plot the average $\operatorname{\mathrm{TCV}}$ across all intermediate layers, for $100$ different random unitary initializations, to avoid slow convergence caused by the depth. As we can see, the $\operatorname{\mathrm{TCV}}$ consistently goes to zero in all runs. Regarding the numerical ranks, we observe that they are all close to the optimal ones ($K=10$).
@@ 5.3 Convergence to $\operatorname{\mathrm{DNC1}}$ Configurations
Figure 3: First three plots from the left: average of different quantities on all intermediate layers and for $100$ different random initializations. Shadows represent the area included between the minimal and maximal values observed. Each optimization step represents 100 epochs. Right: TCV of every layer during training.
@@ Appendix A Useful Definitions and Results
In this section, we will introduce some useful definitions and fact which we will need to prove the results in Appendix˜B. We start recalling the definition and some basic facts about $\Gamma$-convergence of functionals (see (Dal Maso, 1993)).
@@
(Dal Maso, 1993) Let $F_{n}\to F$ in the sense of $\Gamma$-convergence and let $S$ be sequentially compact. Then, for every sequence $x_{n}\in\arg\min_{y\in S}F_{n}(y)$ and every limit point $x$ of $(x_{n})$, we have
@@
The next lemma discusses the zeros of single-variable analytic functions, showing that their cardinality is always finite on compact sets, which we will need later. The results is well known and can be found in (Krantz and Parks, 1992).
Lemma A.3 (Finite number of zero of real analytic functions on compact intervals).
@@
Let $X,Y$ be two finite-dimensional normed spaces, $f:X\times Y\to\mathbb{R}$ and $g:X\times Y\to\mathbb{R}^{m}$ be continuous functions, and $\Omega(x)=\{y\in Y\,|g(x,y)=0\,\}$. Moreover, assume that $f(x,y)$ is coercive with respect to $y$, uniformly in $x$, i.e. that there exists $C>0$ such that $\|f(x,y)\|\geq C\|y\|$ for all $x\in X$. Then, the function
@@
We will use this last result in Section˜B.2 to prove the lower semicontinuity of $F(A,B)=\|AB^{+}\|^{2}$. In fact, by defining
@@
we can define g(x,y)=WBAg(x,y)=WB-A, $f(x,y)=\|W\|^{2}$. Then, recalling that $\Omega(x)=\{y=W\;|\;g(x,y)=WB-A=0\}$, a standard result, see e.g. (Golub and Van Loan, 2013, Section 5.5.1), yields
$F(x)=\inf_{y\in\Omega(x)}f(x,y).$
In particular we are exactly in the setting of Proposition˜A.4, proving lower semicontinuity of $F(A,B)=\|AB^{+}\|^{2}$.
@@
$f(B_{k})$ $=\|AV_{k}^{1}\Sigma_{k,1}^{-1}U_{k,1}^{T}+AV_{k,2}\Sigma_{k,2}^{-1}U_{k,2}^{T}\|\geq\|AV_{k,2}\Sigma_{k,2}^{-1}\|-\|AV_{k,1}\Sigma_{k,1}^{-1}\|$
@@
The main step of the proof is to show that the gradient with respect to a single data point produces matrices of rank at most 1. One then concludes by observing that the sum of $K$ terms of rank at most $1$ has rank at most $K$. Given the structure of the loss function that depends only on the output of the model $f(\mathit{\Theta},x)$, if we use the chain rule and the recursive definition of $f(\mathit{\Theta},x)$ in equation 1, we obtain
@@
$\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big(f_{j}(\mathit{\Theta}_{j},z_{i}),y_{i}\big)=$ $\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big(f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big)$
@@
$R_{i}^{k}(\mathit{\Theta}):=\frac{1}{2}\operatorname{\nabla}_{(z,y)}^{2}\nabla_{W_{l}}\operatorname{\ell}\big(f_{j}(\mathit{\Theta}_{j},\alpha_{i}^{k})\beta_{i}^{k}\big)[(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k}),(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k})],$
for some $(\alpha_{i}^{k},\beta_{i}^{k})$ between $(z_{i},y_{i})$ and $(\bar{z}_{k},\bar{y}_{k})$. Therefore, using the linearity of $\operatorname{\nabla}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}[\cdot]$, we have
@@
$M_{\operatorname{\ell},j}(\mathit{\Theta})$ $=\sup_{k=1,\dots,K}\sup_{i\in\mathcal{C}_{k}}\|\operatorname{\nabla}^{2}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big(f_{j}(\mathit{\Theta}_{j},\alpha_{i}^{k}),\beta_{i}^{k}\big)\|\leq$
$\leq\sup_{k=1,\dots,K}\sup_{(\alpha,\beta)\in\mathcal{H}_{k}^{j}}\|\operatorname{\nabla}^{2}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big(f_{j}(\mathit{\Theta}_{j},\alpha),\beta\big)\|,$
@@
A.3 Proof of Theorem˜3.5
Observe that from equation 4, at the stationary point we have
@@
Now, Lemma 3.4 allows us to control the distance of the gradient in the right-hand side of equation 14 from $\mathcal{M}_{K}^{d_{l-1},d_{l}}$. In particular, given $\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}$, for any $j=1,\dots,l-1$ we have